# countnullcountdimennullnullnulldimennullnullnulldimennullnullnulldimennullnullnulldimen

### v1

@@ Appendix A One-sided concentration under given mean and MAD
$\leq\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\leq 1-\frac{d}{2\mu}.$ (C.1)
The second inequality trivially holds, as we take the infimum over a larger set on the left-hand side compared to the right-hand side. We start by proving the last inequality in (C.1) in Lemma C.1.
Lemma C.1.
It holds that
$\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\leq 1-\frac{d}{2\mu}$
for any $\mu>0$ and $0<d<2\mu$.
Proof.
For any fixed $m$, it is not hard to argue that for some function $f(\mu,d)>0$ independent of $m$, we have $\mathrm{OPT}(m,\mathbb{P}^{m})\geq f(\mu,d)>0$ for every $\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}$, e.g., by using the mechanism that separately sells every good using the optimal robust price in the single good setting. Consider the following case distinction.
@@
$\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\leq\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{f(\mu,d)}=\frac{\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{f(\mu,d)}=0.$ (C.2)
@@
$\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\leq\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{p_{m}}{\mathrm{OPT}(m,\mathbb{P}^{m})}=\frac{p_{m}}{\sup_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\mathrm{OPT}(m,\mathbb{P}^{m})},$ (C.3)
@@
Fix any $\gamma>0$. Using the joint distribution $\mathbb{P}^{m}_{\alpha}=\Pi_{i=1}^{m}\mathbb{P}_{\alpha}$ with $\mathbb{P}_{\alpha}$ the fixed two-point distribution [^6] Any distribution in $\mathcal{P}(\mu,d)$ with finite variance can be used for this argument. with $\alpha=d/(2\mu)$, and using the bundling mechanism that sells at the bundle price $(1-\gamma)m\mu$, gives
$\sup_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\mathrm{OPT}(m,\mathbb{P}^{m})\geq\mathrm{BUND}((1-\gamma)m\mu,\mathbb{P}^{m}_{\alpha})\geq(1-\gamma)m\mu\cdot\left[1-\frac{\mathrm{Var}(X)}{(\gamma\mu)^{2}m}\right]$ (C.4)
by using Lemma A.1. Note that, when $X\sim\mathbb{P}_{\alpha}$,
$\text{Var}(X)=(d/2\mu)\cdot(0-\mu)^{2}+(1-d/(2\mu))\cdot(\mu-\mu-d\mu/(2\mu-d))^{2}=:g(\mu,d).$
Plugging (C.4) into (C.3) gives that for $p_{m}<m(\mu-d/2)$, we have
$\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}$ $\leq\frac{p_{m}}{(1-\gamma)m\mu\cdot\left[1-\frac{g(\mu,d)}{(\epsilon\mu)^{2}m}\right]}\leq\frac{m(\mu-d/2)}{(1-\gamma)m\mu\cdot\left[1-\frac{g(\mu,d)}{(\gamma\mu)^{2}m}\right]}$
$=\frac{1}{1-\gamma}\cdot\frac{2\mu-d}{2\mu}\cdot\left[1-\frac{g(\mu,d)}{(\gamma\mu)^{2}m}\right]^{-1}.$ (C.5)
@@
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\leq\frac{1}{1-\gamma}\cdot\frac{2\mu-d}{2\mu}\cdot\left[1-\frac{g(\mu,d)}{(\gamma\mu)^{2}m}\right]^{-1},$
@@
$\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\leq\frac{1}{1-\gamma}\cdot\frac{2\mu-d}{2\mu}.$
@@
$\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\leq\frac{2\mu-d}{2\mu}.$ (C.6)
@@
We continue with the first inequality in (C.1) in Lemma C.2.
Lemma C.2.
It holds that
$\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\geq\frac{2\mu-d}{2\mu}$
for any $\mu>0$ and $0<d<2\mu$.
@@
For a fixed $0<\epsilon<1-d/(2\mu)$, set
$p_{m}^{*}(\epsilon)=(1-\epsilon)^{2}m\left(\mu-\frac{d}{2(1-\epsilon)}\right)$
for $m\in\mathbb{N}$. We have
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\geq\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m}^{*}(\epsilon),\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\geq\frac{\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\mathrm{BUND}(p_{m}^{*}(\epsilon),\mathbb{P}^{m})}{m\mu}$ (C.7)
using that $\mathrm{OPT}(m,\mathbb{P}^{m})\leq m\mu$. This inequality holds because any truthful mechanism is individually rational, which implies that $\pi(v)\leq\sum_{i}v_{i}$. In expectation, this yields a bound of $m\mu$. For any fixed $\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}$, recall that
$\mathrm{BUND}(p_{m}^{*}(\epsilon),\mathbb{P}^{m})=p_{m}^{*}(\epsilon)\cdot\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq p_{m}^{*}(\epsilon)\right)$ (C.8)
with $p_{m}^{*}(\epsilon)=(1-\epsilon)^{2}m\left(\mu-\frac{d}{2(1-\epsilon)}\right)$.
We next argue that Lemma C.2 follows from Proposition 3.1 by observing that plugging in the concentration bound (3.1) in (C.7) gives
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\geq\frac{(1-\epsilon)^{2}m\left(\mu-\frac{d}{2(1-\epsilon)}\right)(1-f(\mu,d,\epsilon)/m)}{m\mu},$
and then
$\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\geq\frac{(1-\epsilon)^{2}\left(2\mu-\frac{d}{(1-\epsilon)}\right)}{2\mu}.$
Since $\epsilon$ was chosen arbitrarily, the statement of the lemma follows. ∎
This finishes the proof of Theorem 4.1.
Appendix D Proof of Theorem 4.2
First, suppose that the seller chooses $p_{m}^{*}(\epsilon)$ as the selling price. Then
$\inf_{p_{m}}\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}$ $\leq\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m}^{*}(\epsilon),\mathbb{P}^{m})}{m}$
$\leq\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})}{m}$
$-\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m}^{*}(\epsilon),\mathbb{P}^{m})}{m}$
$\leq\mu-\frac{p_{m}^{*}(\epsilon)}{m}(1-f(\mu,d,\epsilon)/m)$
$=\mu-(1-\epsilon)^{2}\left(\mu-\frac{d}{2(1-\epsilon)}\right)(1-f(\mu,d,\epsilon)/m).$
Here we use the upper bound $\mathrm{OPT}(m,\mathbb{P}^{m})\leq m\mu$ and $f(\mu,d,\epsilon)$ is as in Proposition 3.1. Taking the limit of $m\rightarrow\infty$, we then find that
$\lim_{m\rightarrow\infty}\left[\inf_{p_{m}}\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\right]\leq\mu-(1-\epsilon)^{2}\left(\mu-\frac{d}{2(1-\epsilon)}\right).$
Because this bound holds for all $0<\epsilon<1$, by letting $\epsilon\rightarrow 0$ we find that
$\lim_{m\rightarrow\infty}\left[\inf_{p_{m}}\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\right]\leq d/2.$ (D.1)
@@
Case 1: $p_{m}<m(\mu-d/2)$. Note that $\mathrm{BUND}(p_{m},\mathbb{P}^{m})\leq p_{m}<m(\mu-d/2)$ always holds. Also, with a similar argument as in the proof of Lemma C.1, we have for any $0<\gamma<1$, it holds that
@@
This completes the proof.
Appendix E Minimax analysis for $\mathcal{P}_{2}(\mu,d)$ when $\mu\leq d<2\mu$
Proposition E.1.

### v2

@@ 6.1 Competitive ratio under bundling
$\leq\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\leq 1-\frac{d}{2\mu}.$ (6.2)
The second inequality trivially holds, as we take the infimum over a larger set on the left-hand side compared to the right-hand side. We start by proving the first inequality in (6.2).
Consider $p_{m}^{*}$. We have
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\geq\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m}^{*},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\geq\frac{\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\mathrm{BUND}(p_{m}^{*},\mathbb{P}^{m})}{m\mu}$ (6.3)
using that $\mathrm{OPT}(m,\mathbb{P}^{m})\leq m\mu$. This inequality holds because any truthful mechanism is individually rational, which implies that $\pi(v)\leq\sum_{i}v_{i}$. In expectation, this yields a bound of $m\mu$. Then, plugging in the concentration bound (5.2) with $\epsilon=\epsilon_{m}=\frac{1}{2}(1-\frac{d}{2\mu})m^{-1/5}$ in (6.3) gives
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\geq\frac{(1-\epsilon_{m})^{2}m\left(\mu-\frac{d}{2(1-\epsilon_{m})}\right)(1-f(\mu,d)/(m\epsilon_{m}^{4}))}{m\mu},$
and then
$\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\geq\frac{\mu-d/2}{\mu}=1-\frac{d}{2\mu}.$
For any fixed $m$, it is not hard to argue that for some function $g(\mu,d)>0$ independent of $m$, we have $\mathrm{OPT}(m,\mathbb{P}^{m})\geq g(\mu,d)>0$ for every $\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}$, e.g., by using the mechanism that separately sells every good using the optimal robust price in the single good setting. Consider the following case distinction.
@@
$\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\leq\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{g(\mu,d)}=\frac{\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{g(\mu,d)}=0.$ (6.4)
@@
$\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\leq\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{p_{m}}{\mathrm{OPT}(m,\mathbb{P}^{m})}=\frac{p_{m}}{\sup_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\mathrm{OPT}(m,\mathbb{P}^{m})},$ (6.5)
@@
Fix any $\gamma>0$. Using the joint distribution $\mathbb{P}^{m}_{\alpha}=\Pi_{i=1}^{m}\mathbb{P}_{\alpha}$ with $\mathbb{P}_{\alpha}$ the fixed two-point distribution with $\alpha=d/(2\mu)$, and using the bundling mechanism that sells at the bundle price $(1-\gamma)m\mu$, gives
$\sup_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\mathrm{OPT}(m,\mathbb{P}^{m})\geq\mathrm{BUND}((1-\gamma)m\mu,\mathbb{P}^{m}_{\alpha})\geq(1-\gamma)m\mu\cdot\left[1-\frac{\mathrm{Var}(X)}{(\gamma\mu)^{2}m}\right]$ (6.6)
by using Lemma 5.1. Note that, when $X\sim\mathbb{P}_{\alpha}$,
$\text{Var}(X)=(d/2\mu)\cdot(0-\mu)^{2}+(1-d/(2\mu))\cdot(\mu-\mu-d\mu/(2\mu-d))^{2}=:h(\mu,d).$
Plugging (6.6) into (6.5) gives that for $p_{m}<m(\mu-d/2)$, we have
$\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}$ $\leq\frac{p_{m}}{(1-\gamma)m\mu\cdot\left[1-\frac{h(\mu,d)}{(\gamma\mu)^{2}m}\right]}\leq\frac{m(\mu-d/2)}{(1-\gamma)m\mu\cdot\left[1-\frac{h(\mu,d)}{(\gamma\mu)^{2}m}\right]}$
$=\frac{1}{1-\gamma}\cdot\frac{2\mu-d}{2\mu}\cdot\left[1-\frac{h(\mu,d)}{(\gamma\mu)^{2}m}\right]^{-1}.$ (6.7)
@@
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\leq\frac{1}{1-\gamma}\cdot\left(1-\frac{d}{2\mu}\right)\cdot\left[1-\frac{h(\mu,d)}{(\gamma\mu)^{2}m}\right]^{-1},$
@@
$\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\leq\frac{1}{1-\gamma}\cdot\left(1-\frac{d}{2\mu}\right).$
@@
$\lim_{m\rightarrow\infty}\left[\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}_{2}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{\mathrm{OPT}(m,\mathbb{P}^{m})}\right]\leq 1-\frac{d}{2\mu}.$ (6.8)
@@
6.2 Absolute regret under bundling
We now turn to the analysis of absolute regret by proving (3.6).
Proposition 6.2.
Let $0\leq d<2\mu$. The minimax absolute regret satisfies
$\lim_{m\rightarrow\infty}\inf_{p_{m}}\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}=\frac{d}{2},\,$ (6.9)
where the infimum is asymptotically attained by bundling price $p^{*}_{m}$ in (3.3).
@@ 6.2 Absolute regret under bundling
First, suppose that the seller chooses $p_{m}^{*}$ as the selling price. Then
$\inf_{p_{m}}\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\leq\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m}^{*},\mathbb{P}^{m})}{m}$
$\leq\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})}{m}-\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{BUND}(p_{m}^{*},\mathbb{P}^{m})}{m}\leq\mu-\frac{p_{m}^{*}}{m}(1-f(\mu,d)/(m\epsilon_{m}^{4}))$
$=\mu-(1-\epsilon_{m})^{2}\left(\mu-\frac{d}{2(1-\epsilon_{m})}\right)(1-f(\mu,d)/(m\epsilon_{m}^{4})).$
Here we use the upper bound $\mathrm{OPT}(m,\mathbb{P}^{m})\leq m\mu$ and $f(\mu,d)$ is as in Proposition 5.2. Taking the limit of $m\rightarrow\infty$, we then find that
$\lim_{m\rightarrow\infty}\left[\inf_{p_{m}}\sup_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\frac{\mathrm{OPT}(m,\mathbb{P}^{m})-\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\right]\leq\frac{d}{2}.$
@@
Case 1: $p_{m}<m(\mu-d/2)$. Note that $\mathrm{BUND}(p_{m},\mathbb{P}^{m})\leq p_{m}<m(\mu-d/2)$ always holds. Also, we have for any $0<\gamma<1$, it holds that
@@
This completes the proof. ∎
7 Insights from the single-good setting
We now depart from the multi-good setting and consider the single-good case. We characterize the optimal prices under expected revenue, competitive ratio, and absolute regret, and then prove Proposition 3.4.
Optimal prices for expected revenue (Roos et al., 2021) and competitive ratio (Kleer et al., 2024) are given by
$p^{*}_{rev}=\mu\left(1-\sqrt{\frac{d}{2\mu}}\right)\quad\text{and}\quad p^{*}_{cr}=\begin{cases}\mu\left(1-\frac{\sqrt{d(8\mu+d)}-d}{4\mu}\right),\quad&d\leq(2\sqrt{5}-4)\mu,\\ \mu\left(\frac{2\mu-d}{2\mu+d}\right),\quad&d\geq(2\sqrt{5}-4)\mu.\end{cases}$ (7.1)
The optimal price for absolute regret has not previously been characterized in the literature. We therefore derive it from first principles with the full derivation provided in Appendix B. We remark that when the seller is indifferent among prices, we select the smallest one. This tie-breaking rule serves only to preserve continuity of the optimal price without affecting any result or interpretation.
Proposition 7.1.
Let $\mu>0$ and $0\leq d<2\mu$. Then
$p^{*}_{ar}=\arg\inf_{p}\sup_{\mathbb{P}\in\mathcal{P}(\mu,d)}\textup{OPT}(\mathbb{P})-p\mathbb{P}(X\geq p)=\begin{cases}\mu\left(\frac{2\mu-\sqrt{2\mu d}}{2\mu-d}\right),\quad&d\in[0,(3-\sqrt{5})\mu],\\ \mu\left(\frac{d}{2\mu-d}\right),\quad&d\in[(3-\sqrt{5})\mu,\mu],\\ \mu\left(\frac{\mu}{2\mu-d}\right),\quad&d\in[\mu,2\mu).\end{cases}$ (7.2)
We stated earlier that all our results are decreasing in MAD, and that therefore one can replace the exact MAD $d$ by an upper bound without loss of generality. However, this is not true for Proposition 7.1, which we use to prove Proposition 3.4. We therefore derive this result in Appendix B.4 for an upper bound on the MAD as well and show that this does not affect Proposition 3.4.
Let us now consider prices at or above $\mu-d/2$. In this case, Nature can choose a two-point distribution that places all probability mass below the posted price, resulting in zero expected revenue for the seller. Meanwhile, the optimal expected revenue under full information remains positive, albeit small. This outcome renders the competitive ratio equal to zero, even though the outcome is favorable under absolute regret because the absolute revenue loss is small. This clarifies why different objectives can lead to markedly different optimal prices. Moreover, Proposition (7.1) establishes that $p^{*}_{ar}\geq\mu-d/2$ when $d\geq(3-\sqrt{5})\mu$. Combined with the preceding observation, this implies that a seller who optimizes for absolute regret over these MAD values has simultaneously zero expected revenue and a competitive ratio of zero. This explains the appeal of prices that perform well across multiple criteria, such as the bundling price (3.3) in the many-goods setting.
We now establish the proof of Proposition 3.4. First, we show that $p^{*}_{rev}<p^{*}_{cr}$ when $d\leq(2\sqrt{5}-4)\mu$. Hence, we need to show that
$\mu\left(1-\sqrt{\frac{d}{2\mu}}\right)<\mu\left(1-\frac{\sqrt{8d\mu+d^{2}}-d}{4\mu}\right),$
which is equivalent to showing
$4\mu\sqrt{\frac{d}{2\mu}}=\sqrt{8d\mu}>\sqrt{8d\mu+d^{2}}-d.$
This is always true, as $d>0$.
Second, we show that $p^{*}_{rev}<p^{*}_{cr}$ when $d>(2\sqrt{5}-4)\mu$. Hence, we need to show that
$\mu\left(1-\sqrt{\frac{d}{2\mu}}\right)<\mu\left(\frac{2\mu-d}{2\mu+d}\right),$
which is equivalent to showing
$\sqrt{\frac{d}{2\mu}}>\frac{2d}{2\mu+d}.$
Squaring both sides and simplifying leads to $8d\mu<(2\mu+d)^{2}$, which is equivalent to $(2\mu-d)^{2}>0$.
Third, we show that $p^{*}_{cr}<p^{*}_{ar}$ when $d\leq(2\sqrt{5}-4)\mu$. Hence, we need to show that
$\frac{\sqrt{8d\mu+d^{2}}-d}{4\mu}>\frac{\sqrt{2d\mu}-d}{2\mu-d}.$
However, notice that
$\frac{\sqrt{8d\mu+d^{2}}-d}{4\mu}>\frac{\sqrt{8d\mu}-d}{4\mu}=\frac{2\sqrt{2d\mu}-d}{4\mu}.$
Hence, we can show
$\frac{2\sqrt{2d\mu}-d}{4\mu}>\frac{\sqrt{2d\mu}-d}{2\mu-d}.$
instead. This is equivalent to $2\sqrt{2d\mu}<2\mu+d$, which in turn is equivalent to $d<\mu+\frac{d^{2}}{4\mu}$, which is always true, as we assumed $d\leq(2\sqrt{5}-4)\mu<\mu$.
Fourth, we show that $p^{*}_{cr}<p^{*}_{ar}$ when $d\in\left((2\sqrt{5}-4)\mu,\mu\right]$. Hence, we need to show that
$\frac{2\mu-d}{2\mu+d}<\frac{2\mu-\sqrt{2d\mu}}{2\mu-d},$
which is equivalent to showing
$\sqrt{2d\mu}<2\mu-\frac{(2\mu-d)^{2}}{2\mu+d}=\frac{d(6\mu-d)}{2\mu+d}.$
Squaring and division by $d$ results in $2\mu<\frac{d(6\mu-d)^{2}}{(2\mu+d)^{2}}$, which is equivalent to
$h(d):=2\mu(2\mu+d)^{2}-d(6\mu-d)^{2}<0.$
Since $h(d)$ is a third degree polynomial, it has at most three real solutions. One can verify that these solutions are $d_{1}=(6-4\sqrt{2})\mu$, $d_{2}=2\mu$, and $d_{3}=(6+2\sqrt{2})\mu$. Clearly, $d_{1}<(2\sqrt{5}-4)\mu<\mu<d_{2}<d_{3}$ and hence, it suffices to check whether $h(d)<0$ for some $d\in(d_{1},d_{2})$. Choosing $d=\mu$ results in h(d)=7h(d)=-7, which proves our claim. Finally, note that when $d>(3-\sqrt{5})\mu$, then $p^{*}_{ar}>\tau_{1}>p^{*}_{cr}$.
8 Conclusion and future directions
Taken together, our findings indicate that heavy-tailed uncertainty with infinite variance sharply alters the revenue guarantees achievable by any mechanism, eliminating asymptotic full mean extraction. Yet even in this more demanding distributionally robust environment, deterministic bundling remains optimal in the many-goods limit.
We further analyze deterministic bundling through the lenses of competitive ratio and absolute regret. Our results show that these criteria admit a single, robust bundle price when restricting attention to the class of deterministic bundling when $m\xrightarrow{}\infty$. What remains unknown, however, is the speed of this convergence. Quantifying these convergence rates would provide valuable insights into how quickly high-dimensional aggregation smooths out the difference between these three objectives. It also remains unknown whether deterministic bundling is optimal among all truthful mechanisms. Proving this, however, requires a different approach than the one used for the expected revenue objective, since the min-max result for competitive ratio (and the max-min result for absolute regret) are trivial. Any upper bound would therefore be uninformative.
Overall, deterministic bundling remains a powerful tool under heavy-tailed ambiguity, offering strong revenue guarantees and pricing consistency among objectives. Simultaneously, the open questions identified in this work point to rich directions for future research on distributional robustness and high-dimensional mechanism design.
Appendix A Minimax analysis for $\mathcal{P}_{2}(\mu,d)$ when $\mu\leq d<2\mu$
Proposition A.1.